# Effect of generic dark matter halo on transonic accretion onto galactic black holes

### v1

@@ Abstract
The environment surrounding a black hole or black hole binaries is generally expected to play an important role in understanding various astrophysical phenomena around them. In this paper, we study relativistic, low angular momentum, inviscid, and advective hot accretion flow onto a galactic supermassive black hole dressed with a cold dark matter halo. Focusing on different relativistic dark matter distributions with an inner density spike, we analyze the effect of the dark matter halo on the topology and properties of the accretion flow. Our results show enhancement of disk luminosity in the presence of dark matter, which depends on the nature and properties (halo mass and compactness) of the dark matter distribution. Since the dominant contribution to the disk luminosity for compact and massive halo comes from the inner region of the accretion disk, our analysis suggests that luminosity measurement can indeed be useful to probe the exact nature of the dark matter density spike.
@@ I Introduction
The flattening of the rotation curves of spiral and disk galaxies at large radii provides strong evidence of the existence of DM Oort (1940). Studies on the rotation curves of Sgr. A Cautun et al. (2020); de Salas et al. (2019) and other galaxies Rubin et al. (1980); Adams et al. (2014) suggest that over $90\%$ of the galactic mass is in the form of DM. Furthermore, CMB (Cosmic Microwave Background) observations indicate that DM constitutes over $85\%$ of the total energy density of the universe Aghanim et al. (2020). Though the exact nature and origin of DM remain a mystery, its abundance, coldness and lack of strong (non-gravitational) interaction with Standard Model particles are indisputable (see Lisanti (2017) for a review). The cold dark matter (CDM) at the galactic centre gets redistributed by the SMBH, whose strong gravity results in a significant increase (spike) in the local mass density of the CDM outside the event horizon. Assuming the central SMBH (in the absence of DM) to be represented by the Schwarzschild geometry, a fully general relativistic analysis suggests that the DM density distribution must vanish at twice the Schwarzschild radius Sadeghian et al. (2013) (see also Gondolo and Silk (1999) for a Newtonian analysis). Thus, the galactic cores serve as a unique laboratory to probe the effect of DM on strong gravity phenomenology de Laurentis et al. (2023).
@@
Since the dynamics of accretion flow around compact objects is strongly governed by the geometry of the surrounding spacetime, one must accurately model the modification of the spacetime geometry due to the relativistic distribution of the DM particles around the SMBH at the galactic centre. The first step in this scheme is a good description of the galactic DM density profile, guided by observations and large-scale simulations Natarajan (2002). Assuming a Hernquist (HQ) type DM density distribution Hernquist (1990), Cardoso et al.  Cardoso et al. (2022a) proposed an exact analytical solution representing SMBH minimally coupled to the anisotropic dark matter fluid using the Einstein Cluster model Geralico et al. (2012); Einstein (1939) within general relativity (GR). The spacetime solution obtained in  Cardoso et al. (2022a) was used in Cardoso et al. (2022b) to develop a generic, relativistic formalism to study gravitational-wave emission by extreme-mass-ratio inspirals. Going beyond the Hernquist class of DM distribution, a fully numeric pipeline was developed in  Figueiredo et al. (2023); Speeney et al. (2024) to treat generic density perturbation around spherically symmetric galactic SMBH. See Refs. Retana-Montenegro et al. (2012); Jusufi et al. (2020); Igata et al. (2023); Dai et al. (2024); Xavier et al. (2023); Myung (2024); Kazempour et al. (2024); Chen et al. (2024); Tan et al. (2024); Zhao et al. (2024); Mollicone and Destounis (2025); Amancio et al. (2024); Stuchlík and Vrba (2021); Liu et al. (2022); Singha and Biswas (2024); Patra and Majhi (2025); Ranjbar and Olivares-Sánchez (2025); Ranjbar et al. (2025) for studies on the influence of DM halos on different spacetime properties and phenomenology. In the present work, we follow the numerical framework developed in Figueiredo et al. (2023) to model the spacetime around a static spherically symmetric SMBH, dressed with a DM halo, with different relativistic density distributions.
@@
The paper is arranged as follows. In Section II, we discuss the geometry of the spacetime around the central SMBH with different numerically and observationally supported Halo density profiles. In Section III, we highlight the assumptions involved in the analysis and discuss the governing equation for transonic accretion flow. Section IV deals with the occurrence of the sonic points and the topology of the global accretion solution. Subsequently, in Sections V, VI and VII, we discuss the results concerning the parameter space, solution properties and the luminosity of the global accretion solution, respectively. Finally, we conclude with a brief summary and discussion of our results in Section VIII highlighting their importance and prospects.
@@ II.1 Background Geometry
Following Cardoso et al. (2022a); Figueiredo et al. (2023); Speeney et al. (2024); Chakraborty et al. (2024); Pezzella et al. (2024), we employ a generalized Einstein cluster formalism Einstein (1939); Geralico et al. (2012), that considers particles in all possible circular geodesics and provides an average stress tensor that represents an anisotropic fluid with only tangential pressure $P_{t}$. Thus, we assume the metric in Eq. 1 to be a solution of the Einstein’s equations,
@@
is the anisotropic stress-energy tensor encoding the properties of the environment in terms of the density $\rho_{{\rm DM}}(r)$ and the tangential pressure $P_{t}(r)$ of the matter distribution. The anisotropy of the DM distribution is due to the negligible accretion rate of the DM particles onto the central BH. For a given density profile, the continuity equation determines the mass profile as
@@
whereas the metric function $f(r)$ and the tangential pressure $P_{t}(r)$ are determined by the $rr$ component of the field equations (Eq. 2) and the Bianchi identities, respectively,
@@
The stationarity and spherical symmetry of the metric (1) implies the existence of a timelike and a spacelike Killing vector associated with the conserved quantities, the specific energy and the specific angular momentum at infinity, The radius of the innermost stable circular orbit for massive particles is given by the roots of the equation,
@@
To evaluate the metric functions $f(r)$ and $m(r)$, we follow the numerical procedure outlined in Speeney et al. (2024) motivated by the solutions obtained in  Konoplya and Zhidenko (2022); Shen et al. (2024); Maeda et al. (2025). We start with a specific choice of the DM density profile $\rho_{\rm DM}(r)$ (as outlined in Section II.2) and numerically integrate Eq. 4 from $r=2M_{\rm BH}$ to $r_{\infty}=10^{16}M_{\rm BH}$ (corresponding to our numerical infinity) to determine $m(r)$. We then use $m(r)$ in Eq. 5 to evaluate $f(r)$ by integrating backwards from $r_{\infty}$ with the boundary conditions,
@@
where $M_{\rm halo}$ is the total mass of the DM environment surrounding the BH. Note that by construction, the event horizon is $r_{\rm h}=2M_{\rm BH}$. The obtained metric functions can then be used in Eq. 6 to determine the tangential pressure. Similarly, the location and angular momentum at the ISCO and lightring can also be determined explicitly, using Eqs. 7, 8 and 9 respectively.
@@ II.2 Environmental density profiles
where $\rho_{{\rm DM}}^{0}=2^{(\beta-\gamma)/\alpha}\rho_{{\rm DM}}(a_{0})$ is a scale factor with $a_{0}$ being the scale radius of the halo. The parameters $\beta$ and $\gamma$ respectively govern the dependence of the profile at large and small radii, whereas $\alpha$ determines the sharpness of the change of the profile slope at $a_{0}$ Taylor and Silk (2003). The Hernquist and NFW density profiles correspond to the parameter set $(\alpha,\beta,\gamma)$ = $(1,4,1)$ and $(\alpha,\beta,\gamma)$ = $(1,3,1)$, respectively. Since the total mass of the NFW profile is logarithmically divergent, we use a cutoff radius $r_{c}$ for the dark matter distribution, such that $M_{\rm halo}(r>r_{c})=0$. Henceforth, for the NFW density profile, we will use two values of the cutoff radius: (a) $r_{c}=5a_{0}$ (referred to as NFW), (b)$r_{c}=a_{0}$ (referred to as NFW1) . We also use a third-density profile named, Einasto profile Einasto (1969), that, as per recent $N$-body CDM simulations, provides an excellent fit to a wide range of dark matter halos. The Einasto profile is given by,
@@
    We multiply the DM density profile with the cut-off factor $(1-4M_{\rm BH}/r)$. [^1] The radius $r=4M_{\rm BH}$ corresponds to the radius of the unstable circular orbit of a marginally bound particle with angular momentum per unit mass $L=4M_{\rm BH}$. Any particle with energy $\mathcal{E}\leq 1$ and angular momentum $L\geq 4M_{\rm BH}$ has an inner turning point at $r\geq 4M_{\rm BH}$. So, a particle reaching $r=4M_{\rm BH}$ with $\mathcal{E}=1$ is necessarily captured by the BH. The choice of the cutoff factor ensures that the dominant energy condition is satisfied Datta (2024); Speeney et al. (2024) in the region, $2M_{\rm BH}<r<4M_{\rm BH}$.
@@
[…]
@@
    $\begin{split}\bar{\rho}_{{\rm DM}}(r)=&\left(1-\frac{4M_{{\rm BH}}}{r}\right)^{\alpha}\left(\frac{rM_{{\rm BH}}}{M_{{\rm halo}}a_{0}}\right)^{\beta}\\ &\times\left(1+\frac{rM_{{\rm BH}}}{M_{{\rm halo}}a_{0}}\right)^{\gamma},\end{split}$ (15)
@@
Fig. 1 shows the representative plots of different DM distributions for different values of the halo mass and scale radius. We note that the Hernquist and the NFW profiles behave almost similarly, with the Hernquist density being greater than the NFW density at a fixed radius. The NFW distribution terminated at $a_{0}$ (NFW1) has a higher density than the one terminated at $5a_{0}$ since the total halo mass remains constant. For the halo-compactness ($M_{\rm halo}/a_{0}$) considered in Fig. 1, the Hernquist type spike profile has the highest density of DM close to the BH, however for more compact halos, the situation is reversed Speeney et al. (2024).
@@ III Assumptions and Model Equations
In this section, we present the basic equations governing the advection dominated, steady, inviscid, hot accretion flow of non-magnetic and optically thin plasma Yuan and Narayan (2014) onto the central BH. The spherical symmetry of the background geometry (Eq. 1) allows us to consider a planar accretion flow (for brevity we choose to be the equatorial plane). The energy-momentum tensor and the four-current of the fluid is given by,
@@
where $m_{e}$ and $m_{i}$ are the masses of electron and ion, respectively, and $\Theta~{}(=k_{B}T/m_{e}c^{2})$ is the dimensionless temperature. In the above equation and all subsequent calculations, we implicitly assume the number density of the electrons and ions to be the same. Using the REoS in Eq. 20, we express the associated sound-speed as $C_{s}=\sqrt{2\Gamma\Theta/(\tilde{f}+2\Theta)}$, where $\Gamma~{}(=(1+N)/N)$ is the adiabatic index with $N=(1/2)(d\tilde{f}/d\Theta)$ being the polytropic index of the flow. Since the accretion flow is assumed to be confined in the equatorial plane $(\theta=\pi/2,u^{\theta}=0)$, we write the radial component of the relativistic Euler equation as,
@@
As evident from Eq. 23, the effective potential depends on the spacetime geometry and the specific angular momentum of the flow.
The symmetry of the background spacetime allows us to consider stationary and axisymmetric accretion flow. Thus, we define the specific energy ($\mathcal{E}$) and the conserved angular momentum ($\mathcal{L}$) associated with the timelike and spacelike Killing vectors ($\partial_{t}$) and ($\partial_{\phi}$) respectively as,
$-hu_{t}=\mathcal{E},\quad{\rm and}\quad hu_{\phi}=\mathcal{L}~{}.$ (24)
@@
The effect of the DM density profile on the effective potential is shown in Fig. 2.
@@
We observe that compared to the vacuum Schwarzschild BH, the presence of the DM halo lowers the effective potential at any given radius. Thus, a flow with lower specific energy can overcome the potential barrier. For a fixed total halo mass as $a_{0}$ increases (DM distribution gets more dilute), $\Phi_{e}^{\rm eff}$ approaches the Schwarzschild limit for all DM distributions. Again, for a given $a_{0}$, as the halo mass increases, the deviation of the effective potential of each of the DM distributions from the Schwarzschild value increases. However, for a fixed compactness, ($\Psi=M_{\rm halo}/a_{0}$), the higher the halo mass, the lower the effective potential. Though the effective potential for the different DM distributions follows a similar trend, they differ largely among themselves. We observe that close to the central BH, the Einasto profile gives rise to the smallest effective potential. Far away from the central BH the effective potential for the NFW1 profile is usually the smallest. The effective potential for HQ-type DM spike distribution strongly depends on the halo mass as shown in Appendix B.
The radial component of the entropy-generation equation (Eq. 19) gives
@@
The integrated form of the continuity equation (Eq. 16) gives the conserved mass accretion rate of the flow as,
@@
˙$\dot{\cal M}=\frac{\dot{M}}{4\pi\mathcal{K}}=\frac{\rho}{\mathcal{K}}v\gamma_{v}Hr\sqrt{f(r)},$ (27)
where the density is obtained by integrating Eq. 25 using Eqs. 20 and III,
@@
$\mathcal{K}$ being the constant of integration. Since we have considered an adiabatic accretion flow, the entropy accretion rate $\dot{\mathcal{M}}$ remains constant throughout the flow streamline. Furthermore, using Eqs. 25 and 26, we express the dimensionless temperature gradient as,
$\frac{d\Theta}{dr}=$ $-\frac{-2\Theta}{2N+1}$
×$\times\left[\frac{1}{r}\left(\frac{5}{2}+\frac{\lambda^{2}}{r^{2}-\lambda^{2}}\right)+\frac{\gamma_{v}^{2}}{v}\frac{dv}{dr}+\frac{f^{\prime}(r)}{2f(r)}\right].$ (29)
Using Eqs. 22 and III, we obtain the wind equation in terms of the sound speed as,

### v2

@@ Abstract
The environment surrounding a black hole or black hole binaries is generally expected to play an important role in understanding various astrophysical phenomena around them. In this paper, we study relativistic, low angular momentum, inviscid, and advective hot accretion flow onto a galactic supermassive black hole dressed with a cold dark matter halo. Focusing on different relativistic dark matter distributions with an inner density spike, we analyze the effect of the dark matter halo on the topology and properties of the accretion flow. Our results show enhancement of disk luminosity in the presence of dark matter, which depends on the nature and properties (halo mass and compactness) of the dark matter distribution. Under the assumptions of our accretion model, the dominant contribution to the disk luminosity for compact and massive halos arises from the inner regions of the accretion flow. Consequently, our analysis indicates that luminosity measurements can serve as an effective probe of the underlying dark matter density spike.
@@ I Introduction
The flattening of the rotation curves of spiral and disk galaxies at large radii provides strong evidence of the existence of DM Oort (1940). Studies on the rotation curves of Sgr. A Cautun et al. (2020); de Salas et al. (2019) and other galaxies Rubin et al. (1980); Adams et al. (2014) suggest that over $90\%$ of the galactic mass is in the form of DM. Furthermore, CMB (Cosmic Microwave Background) observations indicate that DM constitutes over $85\%$ of the total matter density of the universe Aghanim et al. (2020). Though the exact nature and origin of DM remain a mystery, its abundance, coldness and lack of strong (non-gravitational) interaction with Standard Model particles are indisputable (see Lisanti (2017) for a review). The cold dark matter (CDM) at the galactic centre gets redistributed by the SMBH, whose strong gravity results in a significant increase (spike) in the local mass density of the CDM outside the event horizon. Assuming the central SMBH (in the absence of DM) to be represented by the Schwarzschild geometry, a fully general relativistic analysis suggests that the DM density distribution must vanish at twice the Schwarzschild radius Sadeghian et al. (2013) (see also Gondolo and Silk (1999) for a Newtonian analysis). Thus, the galactic cores serve as a unique laboratory to probe the effect of DM on strong gravity phenomenology de Laurentis et al. (2023).
@@
Since the dynamics of accretion flow around compact objects is strongly governed by the geometry of the surrounding spacetime, one must accurately model the modification of the spacetime geometry due to the relativistic distribution of the DM particles around the SMBH at the galactic centre. The first step in this scheme is a good description of the galactic DM density profile, guided by observations and large-scale simulations Natarajan (2002). Assuming a Hernquist (HQ) type DM density distribution Hernquist (1990), Cardoso et al.  Cardoso et al. (2022a) proposed an exact analytical solution representing SMBH minimally coupled to the anisotropic dark matter fluid using the Einstein Cluster model Geralico et al. (2012); Einstein (1939) within general relativity (GR). The spacetime solution obtained in  Cardoso et al. (2022a) was used in Cardoso et al. (2022b) to develop a generic, relativistic formalism to study gravitational-wave emission by extreme-mass-ratio inspirals. Going beyond the Hernquist class of DM distribution, a fully numeric pipeline was developed in  Figueiredo et al. (2023); Speeney et al. (2024) to treat generic density perturbation around spherically symmetric galactic SMBH. See Refs. Retana-Montenegro et al. (2012); Jusufi et al. (2020); Igata et al. (2023); Dai et al. (2024); Xavier et al. (2023); Myung (2024); Kazempour et al. (2024); Chen et al. (2024); Tan et al. (2024); Zhao et al. (2024); Mollicone and Destounis (2025); Amancio et al. (2024); Stuchl´ık and Vrba (2021); Rahman et al. (2024); Faraji and Rosa (2024); Macedo et al. (2024); Liu et al. (2022); Singha and Biswas (2024); Patra and Majhi (2025); Ranjbar and Olivares-Sánchez (2025); Ranjbar et al. (2025) for studies on the influence of DM halos on different spacetime properties and phenomenology. In the present work, we follow the numerical framework developed in Figueiredo et al. (2023) to model the spacetime around a static spherically symmetric SMBH, dressed with a DM halo, with different relativistic density distributions.
@@
The paper is arranged as follows. In Section˜II, we discuss the geometry of the spacetime around the central SMBH with different numerically and observationally supported Halo density profiles. In Section˜III, we highlight the assumptions involved in the analysis and discuss the governing equation for transonic accretion flow. Section˜IV deals with the occurrence of the sonic points and the topology of the global accretion solution. Subsequently, in Sections˜V, VI and VII, we discuss the results concerning the parameter space, solution properties and the luminosity of the global accretion solution, respectively. Finally, we conclude with a brief summary and discussion of our results in Section˜VIII highlighting their importance and prospects.
@@ II.1 Background Geometry
Following Cardoso et al. (2022a); Figueiredo et al. (2023); Speeney et al. (2024); Chakraborty et al. (2024); Pezzella et al. (2024), we employ a generalized Einstein cluster formalism Einstein (1939); Geralico et al. (2012). Einstein cluster is a collection of collisionless particles in all possible circular geodesics. The average stress energy tensor is given by,
$\left<T^{\mu\nu}\right>=\frac{n}{m_{p}}\left<P^{\mu}P^{\nu}\right$>~,
[…]
@@
is the anisotropic stress-energy tensor encoding the properties of the environment in terms of the density $\rho_{{\rm DM}}(r)$ and the tangential pressure $P_{t}(r)$ of the matter distribution. The zero radial pressure of the distribution is also indicative of the negligible DM accretion rate within the timescale of baryonic accretion. For a given density profile, the continuity equation determines the mass profile as
@@
whereas the metric function $f(r)$ and the tangential pressure $P_{t}(r)$ are determined by the $rr$ component of the field equations (Eq.˜2) and the Bianchi identities, respectively,
@@
The stationarity and spherical symmetry of the metric (1) implies the existence of a timelike and a spacelike Killing vector associated with the conserved quantities, the specific energy and the specific angular momentum at infinity. The radius of the innermost stable circular orbit for massive particles is given by the roots of the equation,
@@
To evaluate the metric functions $f(r)$ and $m(r)$, we follow the numerical procedure outlined in Speeney et al. (2024) motivated by the solutions obtained in  Konoplya and Zhidenko (2022); Shen et al. (2024); Maeda et al. (2025). We start with a specific choice of the DM density profile $\rho_{\rm DM}(r)$ (as outlined in Section˜II.2) and numerically integrate Eq.˜4 from $r=2M_{\rm BH}$ to $r_{\infty}=10^{16}M_{\rm BH}$ (corresponding to our numerical infinity) to determine $m(r)$. We then use $m(r)$ in Eq.˜5 to evaluate $f(r)$ by integrating backwards from $r_{\infty}$ with the boundary conditions,
@@
where $M_{\rm halo}$ is the total mass of the DM environment surrounding the BH. Note that by construction, the event horizon is $r_{\rm h}=2M_{\rm BH}$. The obtained metric functions can then be used in Eq.˜6 to determine the tangential pressure. Similarly, the location and angular momentum at the ISCO and lightring can also be determined explicitly, using Eqs.˜7, 8 and 9 respectively.
@@ II.2 Environmental density profiles
[…]
@@
    We multiply the DM density profile with the cut-off factor $(1-4M_{\rm BH}/r)$. [^1] The radius $r=4M_{\rm BH}$ corresponds to the radius of the unstable circular orbit of a marginally bound particle with angular momentum per unit mass $L=4M_{\rm BH}$. Any particle with energy $\mathcal{E}\leq 1$ and angular momentum $L\geq 4M_{\rm BH}$ has an inner turning point at $r\geq 4M_{\rm BH}$. So, a particle reaching $r=4M_{\rm BH}$ with $\mathcal{E}=1$ is necessarily captured by the BH. The choice of the cutoff factor ensures that the dominant energy condition is satisfied Shen et al. (2025); Datta (2024); Speeney et al. (2024) in the region, $2M_{\rm BH}<r<4M_{\rm BH}$.
@@
    $\begin{split}\rho_{{\rm DM}}(r)=&\bar{\rho}_{{\rm DM}}(r)\frac{M_{{\rm halo}}}{M_{{\rm BH}}^{2}}\\ &\times\left(\int_{4M_{{\rm BH}}}^{r_{c}}4\pi r^{2}\bar{\rho}_{{\rm DM}}(r)dr\right)^{-1},\end{split}$ (14)
@@
    $\bar{\rho}_{{\rm DM}}(r)=$ $\left(1-\frac{4M_{{\rm BH}}}{r}\right)^{\alpha}\left(\frac{rM_{{\rm BH}}}{M_{{\rm halo}}a_{0}}\right)^{\beta}$
    $\times\left(1+\frac{rM_{{\rm BH}}}{M_{{\rm halo}}a_{0}}\right)^{\gamma},$ (15)
@@
Fig. 1 shows the representative plots of different DM distributions for different values of the halo mass and scale radius. We note that the Hernquist and the NFW profiles behave almost similarly, with the Hernquist density being greater than the NFW density at a fixed radius. The NFW distribution terminated at $a_{0}$ (NFW1) has a higher density than the one terminated at $5a_{0}$ since the total halo mass remains constant. For the halo-compactness ($M_{\rm halo}/a_{0}$) considered in Fig.˜1, the Hernquist type spike profile has the highest density of DM close to the BH, however for more compact halos, the situation is reversed Speeney et al. (2024).
@@ III Assumptions and Model Equations
In this section, we present the basic equations governing the advection dominated, steady, inviscid, hot accretion flow of non-magnetic and optically thin plasma Yuan and Narayan (2014) onto the central BH. The spherical symmetry of the background geometry (Eq.˜1) allows us to consider a planar accretion flow (for brevity we choose to be the equatorial plane). The energy-momentum tensor and the four-current of the fluid is given by,
@@
where $m_{e}$ and $m_{i}$ are the masses of electron and ion, respectively, and $\Theta~(=k_{B}T/m_{e}c^{2})$ is the dimensionless temperature. In the above equation and all subsequent calculations, we implicitly assume the number density of the electrons and ions to be the same. Using the REoS in Eq.˜20, we express the associated sound-speed as $C_{s}=\sqrt{2\Gamma\Theta/(\tilde{f}+2\Theta)}$, where $\Gamma~(=(1+N)/N)$ is the adiabatic index with $N=(1/2)(d\tilde{f}/d\Theta)$ being the polytropic index of the flow. Since the accretion flow is assumed to be confined in the equatorial plane $(\theta=\pi/2,u^{\theta}=0)$, we write the radial component of the relativistic Euler equation as,
@@
As evident from Eq.˜23, the effective potential depends on the spacetime geometry and the specific angular momentum of the flow.
The symmetry of the background spacetime allows us to consider a stationary and axisymmetric accretion flow. Thus, we define the specific energy ($\mathcal{E}$) and the conserved angular momentum ($\mathcal{L}$) associated with the timelike and spacelike Killing vectors ($\partial_{t}$) and ($\partial_{\phi}$) respectively as,
$hu_{t}=\mathcal{E},\quad{\rm and}\quad hu_{\phi}=\mathcal{L}~.$ (24)
@@
The effect of the DM density profile on the effective potential is shown in Fig.˜2.
@@
We observe that compared to the vacuum Schwarzschild BH, the presence of the DM halo lowers the effective potential at any given radius. Thus, a flow with lower specific energy can overcome the potential barrier. For a fixed total halo mass as $a_{0}$ increases (DM distribution gets more dilute), $\Phi_{e}^{\rm eff}$ approaches the Schwarzschild limit for all DM distributions. Again, for a given $a_{0}$, as the halo mass increases, the deviation of the effective potential of each of the DM distributions from the Schwarzschild value increases. However, for a fixed compactness, ($\Psi=M_{\rm halo}/a_{0}$), the higher the halo mass, the lower the effective potential. Though the effective potential for the different DM distributions follows a similar trend, they differ largely among themselves. We observe that close to the central BH, the Einasto profile gives rise to the smallest effective potential. Far away from the central BH the effective potential for the NFW1 profile is usually the smallest. The effective potential for HQ-type DM spike distribution strongly depends on the halo mass as shown in Appendix˜B.
The radial component of the entropy-generation equation (Eq.˜19) gives
@@
The integrated form of the continuity equation (Eq.˜16) gives the conserved mass accretion rate of the flow as,
@@
$\dot{\cal M}=\frac{\dot{M}}{4\pi\mathcal{K}}=\frac{\rho}{\mathcal{K}}v\gamma_{v}Hr\sqrt{f(r)},$ (27)
where the density is obtained by integrating Eq.˜25 using Eqs.˜20 and III,
@@
[…]
$\frac{d\Theta}{dr}=$ $-\frac{2\Theta}{2N+1}$
$\times\left[\frac{1}{r}\left(\frac{5}{2}+\frac{\lambda^{2}}{r^{2}-\lambda^{2}}\right)+\frac{\gamma_{v}^{2}}{v}\frac{dv}{dr}+\frac{f^{\prime}(r)}{2f(r)}\right].$ (29)
Using Eqs.˜22 and III, we obtain the wind equation in terms of the sound speed as,